7. Standing on the Earth, you throw an object up in the air. It rises to its highest point, and falls back down to your hand. Which of the following graphs best represents a plot of the vertical velocity vs. time? Circle the correct graph. 7 7 7 7
Answer #1

Upward is +ve and downward is negative

If we throw an object up in air then

Suppose initial velocity is U

then velocity after 't' time will be, during upward mption

V = U + a*t

a = -g, when object is going up, SO

V = U - g*t

it means velocity will decrease with time and at the highest point velocity will be zero.

after that acceleration will be a = +g

during downward motion

V = U + a*t

U = 0 at max height

a = +g

V = g*t

it means with time speed will increase in time,

Velocity will decrease and go to zero then in downward motion (-ve) velocity will increase, which is only followed in Graph 4.

Correct option is graph 4.

Please Upvote.